OpenStudy (anonymous):

A lever with an input arm of 2 meters has a mechanical advantage of 4. What is the output arm's length?

OpenStudy (jamesj):

Use the formula we used in your last question: Advantage = Output/Input. You know two of these variables and hence can calculate the third.

OpenStudy (jamesj):

Here's a short discussion: "Law of the Lever" http://en.wikipedia.org/wiki/Mechanical_advantage#Law_of_the_lever

OpenStudy (anonymous):

@JamesJ Yes, but how do you work it? The mechanical advantage is already given to us. Dont you have to like cross multiply it?

OpenStudy (jamesj):

Yes if Advantage = Output/Input and you know that Advantage = 4 and Input = 2, then what MUST Output be?

OpenStudy (jamesj):

Advantage = Output/Input hence 4 = Output/2 Thus Output = ....?

OpenStudy (anonymous):

@JamesJ I still dont get it

OpenStudy (jamesj):

Let x be the Output length. Then we have 4 = x/2 What MUST the value of x be?

OpenStudy (anonymous):

@JamesJ 2?

OpenStudy (jamesj):

2/2 = 1, so that's not right. And please stop tagging me.

OpenStudy (anonymous):

okay I still dont get it

OpenStudy (jamesj):

What number divided by 2 gives 4. x/2 = 4. What must x be?

OpenStudy (anonymous):

6

OpenStudy (anonymous):

8

OpenStudy (jamesj):

Which one?

OpenStudy (anonymous):

8

OpenStudy (jamesj):

Yes

OpenStudy (jamesj):

4 times 2 is 8

OpenStudy (anonymous):

i knew it but i got confused b/c of cross multiplying

OpenStudy (anonymous):

another one How do you find the input arm? MA = 6 Output arm = .8

OpenStudy (jamesj):

Use EXACTLY the same formula as this problem. MA = Output/Input Now solve for Input

OpenStudy (anonymous):

7.5?

OpenStudy (jamesj):

If MA = Output/Input then Input = Output/MA = 6 / 0.8 = 7.5 Yes

OpenStudy (anonymous):

MA = .5 INPUT ARM = .4 Does OUTPUT ARM = .2?

OpenStudy (jamesj):

MA = Output/Input hence Output = MA * Input = 0.5 * 0.4 = 0.2

OpenStudy (anonymous):

Ramp Problem. A 5-meter ram lifts objects to a height of .75 meters. What is the mechanical advantage of the ramp?